Uranus Company owns 8,000 convertible preference shares of which was acquired in 2020 at cost of P 400,000. The investment was classified as trading securities. On December 31, 2020, the fair value of the preference share was P 425,000. On March 31,2021, Uranus Company converted the 4,000 preference shares into 6,000 shares of ordinary shares, when the market price was P 50 per share for the preferences shares and P 40 per share for the ordinary shares.

00

Based on the above data, answer the following:

1. Compute for the gain on exchange to be recognized in 2021

a. Nil c. P 27,500

b. P 40,000 d. P 60,000

2. The necessary journal entry on march 31, will not include a

a. Debit to Investment in Trading-Ordinary shares, P 240,000

b. Credit to Investment in Trading -Pref, shares, P27,500

c. Credit Gain on exchange, P27,500

d. Credit to Investment in Trading Pref, shares, P 212,500



1
Expert's answer
2021-06-07T11:29:56-0400

Solution:

1.). The correct answer is c. P 27,500.

Gain on exchange is as follows:

The market value is ignored, instead, fair value is used.

First derive the fair value price of the preference shares = "\frac{425,000}{8,000} = 53.125 \; per\; unit" 


Value of 4,000 preference shares converted = 4,000 "\times" 53.125 = 212,500

Value of 6,000 ordinary shares exchanged = 6,000 "\times" 40 = 240,000

Gain on exchange to be recognized in 2021 = 240,000 – 212,500 = P 27,500

 

2.). The correct answer is b. Credit to Investment in Trading -Pref, shares, P27,500


The required journal entries include the following:

                                                                                          P                                   P

Dr. Investment in Trading – Ordinary shares                 240,000

          Cr. Investment in Trading Pref. shares                                                         212,500

          Cr. Gain on exchange                                                                                      27,500

(To record the conversation of 4,000 preference shares to ordinary shares)